Yahaira:

The sum of two numbers is 23. When the second number is subtracted from the first​ number, the difference is 7. Find the two numbers. What are the two​ numbers?

CripQUEZZ:

Ok so 15 and 8 are the only numbers that will work 7 + 8 = 15 15 - 8 = 7 Which means that if you subtract 23 - 15 = You should get 8 23 - 15 = 8

CripQUEZZ:

So your two numbers will be 15 and 8